OpenStudy (anonymous):

Solve the given equation for the slope and y-intercept. x - 6y + 12 = 0 m =? y-intercept =?

OpenStudy (anonymous):

x-6y+12=0 x+12=6y 6y=x+12 y=1/6x+2 y=mx+b m=1/6 b=2

OpenStudy (anonymous):

m is slope b is y intercept